OpenStudy (anonymous):

BUFFERS: Biochemical experiments frequently utilize a buffer system based on TRIS/THAM (HOCH2)3CNH2. The pKa of the conj. acid of TRIS (HOCH2)3CNH3+ is pKa=8.075. What mole ratio ("X") of base-to-acid is required to prepare a buffer at the same pH as human blood, pH=7.4???

OpenStudy (xishem):

Try using the Henderson-Hasselbalch equation...\[pH=pKa+\log(\frac{n_{A^-}}{n_{HA}})\]\[7.4=8.075+\log (\frac{n_{A^-}}{n_{HA}})\]\[-0.675=\log(\frac{n_{A^-}}{n_{HA}})\]\[10^{-0.675}=\ 10^{log(\frac{n_{A^-}}{n_{HA}})}\]\[0.2113=\frac{n_{A^-}}{n_{HA}}\]\[0.211:1.00\]So the ratio of base-to-acid required is... \[1.00:4.73\]